# Blood Flow in Elastic Vessel Using PINNH. Zhang, R. Chan and X.C. Tai

### v1

@@ 4.4 Blood Vessel with Plaque
The initial conditions for the fluid remain the same as in the first problem. The fluid starts at rest at the beginning, as described in Equation (36). The functional $F^{\text{f}}_{\text{init}}$ from Equation (41) remains applicable in this context. The same inlet and outlet boundary conditions are applied as outlined in Equations (37) and (38). However, the interface is changed into $\Gamma^{\text{w}_{1}}\cup\Gamma^{\text{w}_{2}}\cup\Gamma^{\text{p}}$. The continuity of the velocity between the fluid and solid is maintained at the boundaries $\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t}$ by setting the fluid velocity equal to the structure movement speed:
$\mathbf{u}=\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad\text{ on }\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t},t\in[0,T].$ (46)
The equation for the functional $F^{\text{f}}_{\text{bdr}}$, indicating the boundary condition in the fluid problem, is described as follows:
$F^{\text{f}}_{\text{bdr}}=\begin{cases}\boldsymbol{u}-g(0.1-0.1\cos{(2\pi t)})\quad&\text{ in }\Gamma^{\text{ in }}_{t},t\in[0,T],\\ \left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-p\mathbf{I}\right]\cdot\mathbf{n}\quad&\text{ on }\Gamma_{t}^{\text{out}},t\in[0,T],\\ \mathbf{u}-\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad&\text{ on }\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t},t\in[0,T].\end{cases}$ (47)
@@
$\frac{\partial^{2}\eta}{\partial t^{2}}+b_{1}\eta$ $=H_{1}\quad\text{ on }\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{w}_{2}}_{0},$ (48)
@@
$b_{1}=\frac{E_{1}}{\rho_{\text{s}_{1}}(1-\xi^{2})R_{0}^{2}},\quad H=\frac{1}{\rho_{\text{s}_{1}}h_{0}}\left[\frac{R}{R_{0}}P-g\mu\left((\nabla\boldsymbol{u}+(\nabla\boldsymbol{u})^{\intercal})\cdot\boldsymbol{n}\right)\cdot\boldsymbol{e}_{r}\right],$ (49)
@@
$R_{p}(\theta,z)=\begin{cases}R_{0}\quad&\text{ on }\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{w}_{2}}_{0},\\ R_{0}-\sqrt{h_{r}^{2}-\frac{h_{r}^{2}}{h_{l}^{2}}(z-z_{p})^{2}}\quad&\text{ on}\Gamma^{\text{p}}_{0}.\end{cases}$ (50)
For the initial conditions, the fluid remains at rest initially, consistent with (42) in the first problem. Consequently, the functional $F^{\text{s}}_{\text{init}}$ in (44) is directly applicable in this context. Regarding boundary conditions, as the start and end points of the interface boundary $\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{p}}_{0}\cup\Gamma^{\text{w}_{2}}_{0}$ remain the same as those for $\Gamma^{\text{w}}_{0}$, their descriptions do not differ from what is outlined in (43). Thus, we can directly apply the functional (45).
Figure 8: The illustration for the velocity magnitude (up row) and pressure (bottom row) variation at time points 0.01, 0.21, 0.41, 0.61, and 0.81. The red outside indicates the vessel wall. The color inside the lumen indicates the quantity magnitude. For a more convenient observation, the displacement vectors are enlarged by a factor of 10.
@@
The table shows that our method can deliver accurate results, particularly when using $N$, $M$, and $K$ values greater than $1000$. The accuracy of our method surpasses that achieved by using a grid with a minimum length of $\frac{1}{8}\times 10^{-3}$ m and a time step of $1\times 10^{-2}$. This demonstrates the capability of our method for simulating blood flow in complex geometries.
Figure 9: The temporal profiles of the velocity magnitude and pressure at different points. From left to right, the upper three plots illustrate the velocity magnitude on the centerline’s inlet, middle, and outlet. The lower three represent the pressure magnitude for the same corresponding points. The symbols H1, H2, H3, and H4 in the legend denote the grids whose minimum length $h$ are set to be $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$, respectively. Here T1, T2, and T3 denote time step lengths $1\times 10^{-1}$, $1\times 10^{-2}$, and $1\times 10^{-3}$.
@@
Time Step $10^{-1}$ 2.2252 1.7873 0.2809 0.0515 0.0153 $10^{1}$
Time Step $10^{-2}$ 0.2104 0.0776 0.0136 0.0035 0.0094 $10^{2}$
Time Step $10^{-3}$ 0.0885 0.0390 0.0023 0 0.0092 $10^{3}$
@@
Time Step $10^{-1}$ 68.7149 41.6116 0.9274 0.1580 0.0378 $10^{1}$
Time Step $10^{-2}$ 22.8401 7.3519 0.0772 0.0153 0.0188 $10^{2}$
Time Step $10^{-3}$ 0.8589 0.2443 0.0073 0 0.0173 $10^{3}$
Table 4: The comparison between our methods and FEM. The relative errors for velocity magnitude and pressure are reported. The FEM result using a high-resolution grid with a small time step is taken as the ground truth. In the table, $h^{-1}$ signifies the inverse of the minimum length $h$, used to discretize the geometry. Since our method is mesh-free, we test with different numbers of sample points in the column of the results by our method, which is of different time step lengths for FEM.
@@ 4.5 Analysis on Plaque with Different Sizes
Figure 10: The temporal profiles of the stress on the peak point of the plaque (the endpoint of the minor axis of the ellipse) for different plaque sizes.
@@
$F=||\boldsymbol{\sigma}_{\text{f}}\cdot\boldsymbol{n}||_{2}$ (51)
where $\boldsymbol{\sigma}_{\text{f}}$ is the Cauchy stress tensor as in (6), $\boldsymbol{n}$ is the normal vector to the interface between the plaque and fluid. In Figure 10, we present a plot illustrating the force applied by the fluid flow to the highest point of the elliptical plaque. The plot reveals that a healthy vessel without any plaque experiences the least stress in the corresponding region. In contrast, the largest plaque, which significantly alters the vessel geometry, encounters the highest stress from the blood flow. As the minor radius $h_{r}$ of the elliptical plaque increases linearly, the maximum stress within one cardiac cycle rises exponentially. This demonstrates that an increase in the amount of deposited materials within the plaque leads to a much more rapid and significant escalation in the level of risk.
Figure 11: Left: the temporal profiles of the total flux on the outlet section of the vessel for different plaque sizes; Right: total flux within one cardiac cycle measured on the outlet section of the vessel for different plaque sizes.
@@ 5 Conclusion
In this paper, we proposed a mesh-free approach to simulate the blood flow in a deformable vessel. The blood flow is formulated as the Navier-Stokes equation associated with proper initial and boundary conditions. The mechanical model for the vessel wall structure is modeled by equating Newton’s second law of motion and linear elasticity to the force exerted by the fluid flow, which we referred to as the stress continuity equation (6). Our method utilizes the physics-informed neural network (PINN) to solve the Navier-Stokes equation coupled with stress continuity on the interface of vessel structure and fluid flow. As a mesh-free approach, our method does not necessitate discretization and meshing of the computational domain and has been proven highly efficient in solving simulations for complex geometries. With the well-developed neural network packages and parallel modules, our method can be easily accelerated through GPU computing and parallel computing. Furthermore, our model can easily be adapted to arteries in different geometries without requiring additional discretization. Therefore, our method is significantly more efficient and user-friendly than employing the conventional Finite Element Method for clinical purposes. To evaluate our methods, we performed experiments on vessels with and without plaque inside. The results indicate the advantages of our methods. Self-ablation is also performed to determine the optimal setting for network architecture. The time saved by parallel training is evaluated. Additionally, we analyze the influence of different plaque sizes using our method. In future work, we plan to use PINNs to inversely design[20] an optimal shape of stents in patient-specific arteries to help with clinical treatment for patients who encounter arterial stenosis.

### v4

@@ 4.4 Blood Vessel with Plaque
The initial conditions for the fluid remain the same as in the first problem. The fluid starts at rest at the beginning, as described in Equation (32). The function $F^{\text{f}}_{\text{init}}$ from Equation (37) remains applicable in this context. The same inlet and outlet boundary conditions are applied as outlined in Equations (33) and (34). However, the interface is changed into $\Gamma^{\text{w}_{1}}\cup\Gamma^{\text{w}_{2}}\cup\Gamma^{\text{p}}$. The continuity of the velocity between the fluid and solid is maintained at the boundaries $\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t}$ by setting the fluid velocity equal to the structure movement speed:
(42) $\mathbf{u}=\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad\text{ on }\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t},t\in[0,T].$
The equation for the function $F^{\text{f}}_{\text{bdr}}$, indicating the boundary condition in the fluid problem, is described as follows:
(43) $F^{\text{f}}_{\text{bdr}}=\begin{cases}\boldsymbol{u}-g(10-10\cos{(2\pi t)})\quad&\text{ in }\Gamma^{\text{ in }}_{t},t\in[0,T],\\ \left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-P\mathbf{I}\right]\cdot\mathbf{n}\quad&\text{ on }\Gamma_{t}^{\text{out}},t\in[0,T],\\ \mathbf{u}-\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad&\text{ on }\Gamma^{\text{w}_{1}}_{t}\cup\Gamma^{\text{w}_{2}}_{t}\cup\Gamma^{\text{p}}_{t},t\in[0,T].\end{cases}$
@@
(44) $\frac{\partial^{2}\eta}{\partial t^{2}}+b_{1}\eta$ $=H_{1}\quad\text{ on }\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{w}_{2}}_{0},$
@@
(45) $b_{1}=\frac{E_{1}}{\rho_{\text{s}_{1}}(1-\xi^{2})R_{0}^{2}},\quad H=\frac{1}{\rho_{\text{s}_{1}}h_{0}}\left[\frac{R}{R_{0}}P-g\mu\left((\nabla\boldsymbol{u}+(\nabla\boldsymbol{u})^{\intercal})\cdot\boldsymbol{n}\right)\cdot\boldsymbol{e}_{r}\right],$
@@
(46) $R_{p}(\theta,z)=\begin{cases}R_{0}\quad&\text{ on }\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{w}_{2}}_{0},\\ R_{0}-\sqrt{h_{r}^{2}-\frac{h_{r}^{2}}{h_{l}^{2}}(z-z_{p})^{2}}\quad&\text{ on}\Gamma^{\text{p}}_{0}.\end{cases}$
For the initial conditions, the fluid remains at rest initially, consistent with (38) in the first problem. Consequently, the function $F^{\text{s}}_{\text{init}}$ in (40) is directly applicable in this context. Regarding boundary conditions, as the start and end points of the interface boundary $\Gamma^{\text{w}_{1}}_{0}\cup\Gamma^{\text{p}}_{0}\cup\Gamma^{\text{w}_{2}}_{0}$ remain the same as those for $\Gamma^{\text{w}}_{0}$, their descriptions do not differ from what is outlined in (39). Thus, we can directly apply the function (41).
Figure 8: Illustration of velocity magnitude (first row), pressure (second row), displacement magnitude (third row) and displacement magnitude for wall (forth row) variations at time points 0.01, 0.21, 0.41, 0.61, and 0.81. The first and the second don’t contain the vessel wall. Deformations of the shape are enlarged by a factor of 10 for better visibility.
@@
The table shows that our method can deliver accurate results, particularly when using $N$, $M$, and $K$ values greater than $1000$. The accuracy of our method surpasses that achieved by using a grid with a minimum length of $\frac{1}{8}\times 10^{-3}$ meter and a time step of $1\times 10^{-2}$ seconds. This demonstrates the capability of our method for simulating blood flow in complex geometries.
Figure 9: The temporal profiles of the velocity magnitude and pressure at different points. From left to right, the upper three plots illustrate the velocity magnitude on the centerline’s inlet, middle, and outlet. The lower three represent the pressure magnitude for the same corresponding points. The symbols H1, H2, H3, and H4 in the legend denote the grids whose minimum length $h$ are set to be $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$ meter, respectively. Here T1, T2, and T3 denote time step lengths $1\times 10^{-1}$, $1\times 10^{-2}$, and $1\times 10^{-3}$ second.
@@
Time Step $10^{-1}$ 2.2252 1.7873 0.2809 0.0515 0.0153 $10^{2}$
Time Step $10^{-2}$ 0.2104 0.0776 0.0136 0.0035 0.0094 $10^{3}$
Time Step $10^{-3}$ 0.0885 0.0390 0.0023 0 0.0092 $10^{4}$
@@
Time Step $10^{-1}$ 68.7149 41.6116 0.9274 0.1580 0.0378 $10^{2}$
Time Step $10^{-2}$ 22.8401 7.3519 0.0772 0.0153 0.0188 $10^{3}$
Time Step $10^{-3}$ 0.8589 0.2443 0.0073 0 0.0173 $10^{4}$
Computation Time (seconds)
$h^{-1}$ $2\times 10^{3}$ $4\times 10^{3}$ $8\times 10^{3}$ $16\times 10^{3}$ Ours Points
Time Step $10^{-1}$ 53 147 411 3778 2832 $10^{2}$
Time Step $10^{-2}$ 241 527 4671 20512 3163 $10^{3}$
Time Step $10^{-3}$ 761 3274 18375 147381 9417 $10^{4}$
Table 5: Comparison between our methods and FEM on cylinder vessel with plaques. The FEM result on a high-resolution grid with a small time step is considered ground truth. $h^{-1}$ signifies the inverse of the minimum length $h$, used to discretize the geometry. Results by our method are shown with different numbers of sample points, corresponding to different time step lengths for FEM.
@@ 4.5 Analysis on Plaque with Different Sizes
(47) $F=||\boldsymbol{\sigma}_{\text{f}}\cdot\boldsymbol{n}||_{2}$
where $\boldsymbol{\sigma}_{\text{f}}$ is the Cauchy stress tensor as in (6), $\boldsymbol{n}$ is the normal vector to the interface between the plaque and fluid. In the left plot of Figure 10, we present a plot illustrating the force applied by the fluid flow to the highest point of the elliptical plaque. The plot reveals that a healthy vessel without any plaque experiences the least stress in the corresponding region. In contrast, the largest plaque, which significantly alters the vessel geometry, encounters the highest stress from the blood flow. As the minor radius $h_{r}$ of the elliptical plaque increases linearly, the maximum stress within one cardiac cycle rises exponentially. This demonstrates that an increase in the amount of deposited materials within the plaque leads to a much more rapid and significant escalation in the level of risk.
Figure 10: Left: the temporal profiles of the stress on the peak point of the plaque for different plaque sizes; Middle: the temporal profiles of the total flux on the outlet section of the vessel for different plaque sizes; Right: total flux within one cardiac cycle measured on the outlet section of the vessel for different plaque sizes.
@@ 5 Conclusion
In this paper, we proposed a mesh-free approach to simulate the blood flow in a deformable vessel. The blood flow is formulated as the Navier-Stokes equation associated with proper initial and boundary conditions. The mechanical model for the vessel wall structure is modeled by equating Newton’s second law of motion and linear elasticity to the force exerted by the fluid flow, which we referred to as the stress continuity equation (6). Our method utilizes the physics-informed neural network (PINN) to solve the Navier-Stokes equation coupled with stress continuity on the interface of vessel structure and fluid flow. As a mesh-free approach, our method does not necessitate discretization and meshing of the computational domain and has been proven highly efficient in solving simulations for complex geometries. With the well-developed neural network packages and parallel modules, our method can be easily accelerated through GPU computing and parallel computing. Furthermore, our model can easily be adapted to arteries in different geometries without requiring additional discretization. Therefore, our method is significantly more efficient and user-friendly than employing the conventional Finite Element Method for clinical purposes. To evaluate our methods, we performed experiments on vessels with and without plaque inside. The results indicate the advantages of our methods. Self-ablation is also performed to determine the optimal setting for network architecture. The time saved by parallel training is evaluated. Additionally, we analyze the influence of different plaque sizes using our method.
In our future work, we aim to leverage Physics-Informed Neural Networks (PINNs) for the inverse design of optimal stent shapes in patient-specific arteries, aiding in the clinical treatment of arterial stenosis [23]. The interaction to non-linear elastic and porous solid structure are also important to model. Additionally, given the current capabilities of imaging techniques to capture precise and high-resolution cardiovascular vessel deformations, investigating flow simulation within dynamic domains presents an intriguing and practical avenue for research.
When considering network architectures for our PINN models, we plan to explore more advanced and innovative designs. For instance, integrating principles of domain decomposition, as demonstrated in approaches like xPINN [32, 17], could enhance the accuracy of our results, particularly when addressing challenges beyond the capabilities of other methods. Furthermore, exploring architectures that employ adaptive activation functions [18] instead of the alternating Sigmoid and ReLU setup could potentially lead to improved convergence rates and overall performance.
Appendix A One-Pulse Flow Simulation
Figure 11: Illustration of velocity magnitude (first row), pressure (second row), displacement magnitude (third row) and displacement magnitude for wall (forth row) variations at time points 0.04, 0.07, 0.10, and 0.14. The first and the second don’t contain the vessel wall. Deformations of the shape are enlarged by a factor of 10 for better visibility.
Here we perform an experiment with an impulse input at the inlet of the vessel as the inlet boundary condition so that a dilation phenomenon that propagates downward to the outlet can observed. Within this experiment, the fluid density $\rho$ is set to $1.025$ g/$\text{cm}^{3}$. The viscosity $\mu$ is $0.035$ poise. The vessel has a cylinder-like shape, with a length $l$ of $25$ cm and a radius $r_{0}$ of $1$ cm. The vessel wall thickness is $h=0.2$ cm. The vessel structure density is set as $\rho_{\text{s}}=1.2$ g/$\text{cm}^{3}$, with Young modulus as $E_{\text{s}}=0.8\times 10^{7}$ dynes/$\text{cm}^{2}$, and a Poisson’s ratio of $\xi_{\text{s}}=0.5$. The entire simulation runs for a duration $T$ of $0.2$ seconds.
Fluid enters the vessel through the inlet, following a parabolic velocity profile $g(\boldsymbol{x})=1-\frac{r^{2}}{r_{0}^{2}}$. Equation (48) provides the mathematical representation for this inlet velocity profile. The impulse input increases and decreases very quickly so that the flow will build up a segment of high pressure part that flows down. In our setting, the velocity speed at the input will finish increasing and decreasing within 0.1 second, which is described as:
(48) $\boldsymbol{u}(\boldsymbol{x},t)=g(\boldsymbol{x})(10-10\cos{(20\pi t)})\quad\text{ on }\Gamma^{\text{ in }}_{t},t\in[0,T].$
Other settings like outlet boundary conditions are kept as same as those in the previous cylinder vessel simulation problem. The solved result is visualized in Figure.11, whose layout and color visualized follow the same rule as those in the main context. The frames rendered from left to right are at times 0.04, 0.07, 0.10, and 0.14. The dilation propagation can be clearly observed.
Appendix B Navier-Stokes Equation in Cylindrical System
In our problem, since the fluid is incompressible, we can have $\boldsymbol{\sigma}_{f}=P\boldsymbol{I}+2\mu\nabla\boldsymbol{u}$ as $\nabla\cdot(\nabla\boldsymbol{u})^{T}=0$ according to [29, P.31, Sec. 8]. Then, the Navier-Stokes equation in polar coordinates takes the following form
(49) $\rho_{\text{f}}\frac{\partial u_{z}}{\partial t}+\rho_{\text{f}}(u_{r}\frac{\partial}{\partial r}+\frac{u_{\theta}}{r}\frac{\partial}{\partial\theta}+u_{z}\frac{\partial}{\partial z})u_{z}$ $=-\frac{\partial P}{\partial z}+\mu\nabla^{2}u_{z}$
$\rho_{\text{f}}\frac{\partial u_{r}}{\partial t}+\rho_{\text{f}}(u_{r}\frac{\partial}{\partial r}+\frac{u_{\theta}}{r}\frac{\partial}{\partial\theta}+u_{z}\frac{\partial}{\partial z})u_{r}-\frac{u_{\theta}^{2}}{r}$ $=-\frac{\partial P}{\partial r}+\mu\left(\nabla^{2}u_{r}-\frac{u_{r}}{r^{2}}-\frac{2}{r^{2}}\frac{\partial u_{\theta}}{\partial\theta}\right)$
$\rho_{\text{f}}\frac{\partial u_{\theta}}{\partial t}+\rho_{\text{f}}(u_{r}\frac{\partial}{\partial r}+\frac{u_{\theta}}{r}\frac{\partial}{\partial\theta}+u_{z}\frac{\partial}{\partial z})u_{\theta}+\frac{u_{r}u_{\theta}}{r}$ $=-\frac{1}{r}\frac{\partial P}{\partial\theta}+\mu\left(\nabla^{2}u_{\theta}+\frac{u_{\theta}}{r^{2}}+\frac{2}{r^{2}}\frac{\partial u_{r}}{\partial\theta}\right)$
where $\nabla^{2}=\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}}{\partial\theta^{2}}+\frac{\partial^{2}}{\partial z^{2}}$. The incompressibility reads,
(50) $\frac{1}{r}\frac{\partial}{\partial r}(ru_{r})+\frac{1}{r}\frac{\partial u_{\theta}}{\partial\theta}+\frac{\partial u_{z}}{\partial z}=0$
and the harmonic extension is as:
(51) $\frac{1}{r}\frac{\partial}{\partial r}(r\frac{\partial u_{z}}{\partial r})+\frac{1}{r^{2}}\frac{\partial^{2}u_{z}}{\partial\theta^{2}}+\frac{\partial^{2}u_{z}}{\partial z^{2}}=0$
$\frac{1}{r}\frac{\partial}{\partial r}(r\frac{\partial u_{r}}{\partial r})+\frac{1}{r^{2}}\frac{\partial^{2}u_{r}}{\partial\theta^{2}}+\frac{\partial^{2}u_{r}}{\partial z^{2}}=0$
$\frac{1}{r}\frac{\partial}{\partial r}(r\frac{\partial u_{\theta}}{\partial r})+\frac{1}{r^{2}}\frac{\partial^{2}u_{\theta}}{\partial\theta^{2}}+\frac{\partial^{2}u_{\theta}}{\partial z^{2}}=0$
As we assume that our problem is axis-symmetric, we have $u_{\theta}=0$ for all $\theta$. Accordingly, (49) is reduced into
(52) $\rho_{\text{f}}\frac{\partial u_{z}}{\partial t}+\rho_{\text{f}}(u_{r}\frac{\partial}{\partial r}+u_{z}\frac{\partial}{\partial z})u_{z}$ $=-\frac{\partial P}{\partial z}+\mu\left(\frac{1}{r}\frac{\partial u_{z}}{\partial r}+\frac{\partial^{2}u_{z}}{\partial r^{2}}+\frac{\partial^{2}u_{z}}{\partial z^{2}}\right)$
$\rho_{\text{f}}\frac{\partial u_{r}}{\partial t}+\rho_{\text{f}}(u_{r}\frac{\partial}{\partial r}+u_{z}\frac{\partial}{\partial z})u_{r}$ $=-\frac{\partial P}{\partial r}+\mu\left(\frac{1}{r}\frac{\partial u_{r}}{\partial r}+\frac{\partial^{2}u_{r}}{\partial r^{2}}+\frac{\partial^{2}u_{r}}{\partial z^{2}}-\frac{u_{r}}{r^{2}}\right)$
and the incompressibility (50) is reduced into
(53) $\frac{1}{r}u_{r}+\frac{\partial u_{r}}{\partial r}+\frac{\partial u_{z}}{\partial z}=0$
and the harmonic extension (51) is reduced into
(54) $\frac{1}{r}\frac{\partial u_{z}}{\partial r}+\frac{\partial^{2}u_{z}}{\partial r^{2}}+\frac{\partial^{2}u_{z}}{\partial z^{2}}=0$
$\frac{1}{r}\frac{\partial u_{r}}{\partial r}+\frac{\partial^{2}u_{r}}{\partial r^{2}}+\frac{\partial^{2}u_{r}}{\partial z^{2}}=0$
The cylindrical form of a PDE would introduce $\frac{1}{r}$ which results in singularities at $r=0$. To deal with it, we first map $[0,r_{0}]\times[0,2\pi)$ into $[-r_{0},r_{0}]\times[0,\pi)$ and let $r^{\prime}=sgn(r)(ReLU(|r|-\epsilon_{r})+\epsilon_{r})$ so that $|\frac{1}{r^{\prime}}|$ is bounde by $\frac{1}{\epsilon_{r}}$ near $r=0$, where $sgn(r)$ denotes the sign of $r$. Then the residual form in the cylindrical representation for Navier-Stokes equation is as
(55) $\mathcal{L}_{\text{ns}}({{\boldsymbol{\theta}_{f}}},{{\boldsymbol{\theta}_{d}}})$
$=||\rho_{\text{f}}\frac{\partial u_{z}^{*}}{\partial t}+\rho_{\text{f}}(u_{r}^{*}\frac{\partial}{\partial r}+u_{z}^{*}\frac{\partial}{\partial z})u_{z}^{*}+\frac{\partial P^{*}}{\partial z}-\mu\left(\frac{1}{r^{\prime}}\frac{\partial u_{z}^{*}}{\partial r}+\frac{\partial^{2}u_{z}^{*}}{\partial r^{2}}+\frac{\partial^{2}u_{z}^{*}}{\partial z^{2}}\right)||^{2}_{L^{2}(\Omega_{t}^{\text{f}};[0,T])}$
$+||\rho_{\text{f}}\frac{\partial u_{r}^{*}}{\partial t}+\rho_{\text{f}}(u_{r}^{*}\frac{\partial}{\partial r}+u_{z}^{*}\frac{\partial}{\partial z})u_{r}^{*}+\frac{\partial P^{*}}{\partial r}-\mu\left(\frac{1}{r^{\prime}}\frac{\partial u_{r}^{*}}{\partial r}+\frac{\partial^{2}u_{r}^{*}}{\partial r^{2}}+\frac{\partial^{2}u_{r}^{*}}{\partial z^{2}}-\frac{u_{r}^{*}}{r^{\prime 2}}\right)||^{2}_{L^{2}(\Omega_{t}^{\text{f}};[0,T])}$
$+||\frac{1}{r^{\prime}}u_{r}^{*}+\frac{\partial u_{r}^{*}}{\partial r}+\frac{\partial u_{z}^{*}}{\partial z}||^{2}_{L^{2}(\Omega_{t}^{\text{f}};[0,T])}$
where $(u_{z}^{*},u_{r}^{*})=\boldsymbol{u}^{*}$ and $P^{*}$ are calculated from (21), whose input $\boldsymbol{x}_{t}^{*}$ is calculated from (20) by the input $\boldsymbol{x}_{0}=(r,z)$. We set $\epsilon_{r}=\frac{r_{0}}{100}$ in the implementation.
Appendix C Mini-batch gradient descent algorithm
(56) $\mathcal{S}({\Omega_{0}^{\text{f}}};t=0)=\{(\boldsymbol{x}^{(i)}_{0},0)\quad|\quad\text{ where }\boldsymbol{x}^{(i)}_{0}\in\mathcal{S}({\Omega_{0}^{\text{f}}})\}$
(57) $\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])=\{(\boldsymbol{x}^{(i)}_{t^{(i)}},t^{(i)})\quad|$ $\boldsymbol{x}^{(i)}_{t^{(i)}}:=\boldsymbol{x}^{(i)}_{0}+\boldsymbol{d}^{*}(\boldsymbol{x}^{(i)}_{0},t^{(i)};\boldsymbol{\theta}_{d})$
[…]
(58) $||f||^{2}_{\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])}:=\frac{1}{M}\sum\limits_{(\boldsymbol{x},t)\in\mathcal{S}({\Omega_{t}^{\text{f}}};[0,T])}|f(\boldsymbol{x},t)|^{2}$
(59) $||\cdot||^{2}_{L^{2}(\cdot;\cdot)}:=||\cdot||^{2}_{\mathcal{S}(\cdot;\cdot)}$
Appendix D Parallel Training
(60) $\nabla_{{\boldsymbol{\theta}_{d}}}\mathcal{L}_{\text{solid}}(\boldsymbol{\theta}_{\text{d}};\mathcal{N})=\frac{1}{P}\sum\limits_{i=1}^{P}\nabla_{{\boldsymbol{\theta}_{\text{d}}^{(i)}}}\mathcal{L}_{\text{solid}}(\boldsymbol{\theta}_{\text{d}}^{(i)};\mathcal{N}^{(i)})$